OpenStudy (anonymous):

Eliana and her sister are comparing the prices of 2 brans of cereals. Toasty oats costs $2.25 fir a 15-ounce box. Crunchy Oaties costs $390 for a 30-ounce bag. A. What is the price per ounce of the Toasty Oats? B. How much more expensive is Toasty Oats per ounce than Crunchy Oaties

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@LoveYou*69

OpenStudy (anonymous):

please help

hero (hero):

@LoveYou*69 can actually help you? I'd like to see it

hero (hero):

I'd also like a box of Crunchy Oaties

OpenStudy (anonymous):

ugh please just help i have 11 more questions to finish after this

OpenStudy (anonymous):

A) 15¢ B) 2¢ more I think, hero you can tell me if im wrong... thats 3.90 not 390 dollars right? lol

hero (hero):

I'm pretty sure you mean $3.90 and not $390

OpenStudy (anonymous):

but i need to show work! how do you get that

hero (hero):

Hint: \[\frac{2.25 \div 15 }{15 \div 15} = \frac{.15}{1}\]

OpenStudy (anonymous):

A) 2.25 / 15 = 0.15 B) 3.90/ 30 = 0.13 so its 2¢ LESS (not more lol i messed up)

hero (hero):

Second Hint 3.90 - 2.25 = 1.65

OpenStudy (anonymous):

so is b /13 or 1.65?

hero (hero):

If you can't figure it out from there, I can't help you

OpenStudy (anonymous):

.13*

hero (hero):

Hint: b is not .13 but nice try @LoveYou*69

OpenStudy (anonymous):

im not its my homework

OpenStudy (anonymous):

wait so is b 1.65?

OpenStudy (anonymous):

-_- hero so its not 13¢ per ounce for the second box? which would mean QB is 2¢ cause thats how much more the first box per ounce is compared to the second?

OpenStudy (anonymous):

@Hero